OpenStudy (anonymous):

I was hoping to get help with verifying/establishing this identity: 4sin^2 theta+ 2cos^ theta= 4- 2cos^ theta

zepdrix (zepdrix):

\[\Large\rm 4(\color{red}{\sin^2 \theta})+2\cos^2\theta\]Use your Pythagorean Identity on the red part.

zepdrix (zepdrix):

I put brackets around it so you'll remember to `distribute` the 4 to each term that shows up. Can be easy to forget otherwise.

zepdrix (zepdrix):

You remember your identity, yes? :) \[\Large\rm \sin^2\theta+\cos^2\theta=1\qquad\to\qquad \color{red}{\sin^2\theta=1-\cos^2\theta}\]

OpenStudy (anonymous):

Yes, I remember the identity. Thank you so much again for demonstrating it for me!